PH of a buffer solution: Mass of NaC2H3O2 * 3H2O (FW=136g/mol): 3.540g Calculate the Theoretical ...

pH of a buffer solution: Mass of NaC2H3O2 * 3H2O (FW=136g/mol): 3.540g

Calculate the Theoretical pH of these buffer solutions:

A. pH of original buffer solution (with 8.8 mL of 3.0 M acetic acid and 55.6 mL of water) : 4.7

B. pH of buffer + 1.0 mL of 6.0 M HCl: 3.8

C. pH of buffer + 1.0 mL of 6.0 M NaOH: 4.9
Answer #1

3.540 g NaC2H3O2 * 3H2O (salt) = mass / molar mass = 3.540 g / 136 g / mole = 0.026 mole.

8.8 mL of 3.0 M acetic acid = 0.0088 L * 3.0 mole / L = 0.0264 mole.

a) Using Henderson equation,

pH = pKa + log [salt] / [acid]

or

pH = 4.74 + log (0.026 / 0.0264)

or

pH = 4.73

b) 1.0 mL of 6.0 M HCl = 0.001 L * 6.0 mole / L = 0.006 mole.

pH = pKa + log [salt] / [acid]

or

pH = 4.74 + log [(0.026 - 0.006) / (0.0264 + 0.006)]

or

pH = 4.53

c) 1.0 mL of 6.0 M NaOH = 0.001 L * 6.0 mole / L = 0.006 mole.

pH = pKa + log [salt] / [acid]

or

pH = 4.74 + log [(0.026 + 0.006) / (0.0264 - 0.006)]

or

pH = 4.93